Circular No. 7182

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999cj AND 1999ck IN ANONYMOUS GALAXIES D. Hardin, Laboratoire de Physique Nucléaire et des Hautes Énergies,

CNRS/IN2P3, and Universités Paris VI et VII, on behalf of the European Supernova Cosmology Consortium (N. A. Walton, R. S. Ellis, M. Irwin, R. G. McMahon, I. Hook, P. Ruiz-Lapuente, J. Mendez, A. Visco, P. Astier, C. Balland, G. Blanc, A. Blanchard, S. Fabbro, F. Hammer, D. Hardin, A. Letessier-Selvon, J. M. Lévy, M. Mouchet, R. Pain, J. Rich, K. Shahmanèche, R. Taillet), and in collaboration with the Supernova Cosmology Project (cf. IAUC 7125, 7128), report two type-Ia supernovae found with the Isaac Newton 2.5-m telescope in their 1999 Intermediate Redshift Supernova Search:

SN  1999 UT  α_{2000}  δ_{2000}  R  Offset
1999cj Apr. 17 10^{h}14^{m}02.^{s}46 − 0^{o}14′02′′.5 19.2 2′′.5 E, 0′′.6 S
1999ck Apr. 16 14 08 56.59 − 0 05 50.7 21.9 1′′.0 E, 0′′.8 S

SN1999cj was discovered ∼ 7 days past maximum. SN1999ck was discovered ∼ 5 days before maximum. Both objects were confirmed with spectra obtained on Apr. 19 at the William Herschel Telescope.

COMET C/1999 K8 (LINEAR) Another apparently asteroidal object discovered by LINEAR (discov-

ery observation below; additional astrometry on MPEC 1999-K44), and posted on The NEO Confirmation Page, has been reported as cometary on May 27 by L. Šarounová (Ondřejov; coma diameter at least 30′′ with condensed nucleus), by M. Tichý and Z. Moravec (Kleť; 10′′ coma), and by R. A. Koff (Thornton, CO; diffuse coma of diameter ∼ 8′′).

1999 UT  α_{2000}  δ_{2000}  m_{2}
May 26.38431  21^{h}35^{m}58.^{s}89  +34^{o}29′03′′.0  18.6

Very preliminary parabolic orbital elements by B. G. Marsden:

T = 2000 Jan. 17.585 TT  ω = 144.^{o}350
Ω = 199.122 }2000.0
q = 4.84481 AU  i = 51.540
1999 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green